A price-weighted index consists of stocks A, B, and C which are priced at $55, $38, and $17 a share, respectively. The current index divisor is 2.5. What will the new index divisor be if stock A undergoes a 5-for-1 stock split?
Answer #1

Before stock split

Stock Price
A 55
B 38
C 17

Current index divisor = 2.5

Price weighted Index = Sum of prices of the stocks/Current Index number = (55+38+17)/2.5 = 110/2.5 = 44

After stock split

Stock A undergoes 5-for-1 stock split i.e., one stock is split into 5 stocks

The new price of A = 55/5 = 11

Following table shows the prices of stocks after the stock split of A

Stock Price
A 11
B 38
C 17

After the stock, the price-weighted index will not change and only the index divisor will change

New Index divisor = n

New price-weighted index = (11+38+17)/n = 44

66/n = 44

n = 66/44 = 1.5

Answer

New Index divisor (after stock split of A) = 1.5